# Dual Diffusion for Unified Image Generation and Understanding

### v1

@@ Abstract
Diffusion models have gained tremendous success in text-to-image generation, yet still lag behind with visual understanding tasks, an area dominated by autoregressive vision-language models. We propose a large-scale and fully end-to-end diffusion model for multi-modal understanding and generation that significantly improves on existing diffusion-based multimodal models, and is the first of its kind to support the full suite of vision-language modeling capabilities. Inspired by the multimodal diffusion transformer (MM-DiT) and recent advances in discrete diffusion language modeling, we leverage a cross-modal maximum likelihood estimation framework that simultaneously trains the conditional likelihoods of both images and text jointly under a single loss function, which is back-propagated through both branches of the diffusion transformer. The resulting model is highly flexible and capable of a wide range of tasks including image generation, captioning, and visual question answering. Our model attained competitive performance compared to recent unified image understanding and generation models, demonstrating the potential of multimodal diffusion modeling as a promising alternative to autoregressive next-token prediction models.
@@ 2.1 Continuous Diffusion
where velocity field $\mathbf{v}(\mathbf{x}_{t},t)=\dot{\alpha_{t}}\mathbf{x}+\dot{\sigma_{t}}\mathbf{\epsilon}$. The ODE in (2) pushes the distribution of $\mathbf{x}_{t}$ from $p_{0}$ to $p_{T}$. To generate new samples, we can use neural networks to approximate the velocity field $\mathbf{v}$ and then integrate ODE (2) backward in time starting from $\mathbf{x}_{T}\sim\mathcal{N}(0,I)$. A common choice of $\alpha_{t},\sigma_{t}$ in flow matching model is $\alpha_{t}=1-t,\sigma=t$ and therefore $\mathbf{v}=\epsilon-\mathbf{x}$, which corresponds to the optimal transport interpolant between two distribution $p_{0}$ and $p_{1}$ [45, 50]. The neural network for regressing the velocity field $\mathbf{v}$ in (2) is trained by optimizing the flow matching loss
@@ 2.2 Discrete Diffusion
In discrete diffusion, the variate $\mathbf{x}\in\mathcal{X}\times\dots\times\mathcal{X}$ has finite support over the product space of $\mathcal{X}=\{1,\dots,N\}$, where in language models $N$ is the vocabulary size of the token embedding. Generally, there are two ways to approach this modeling task. The first line of works [41, 12, 15, 52, 27] apply a continuous relaxation to the discrete variable and proceed with a continuous reformulation of the framework, allowing the application of the equations in Section 2.1. This greatly simplifies the diffusion modeling itself, but introduces a significant source of error in the mapping between discrete and relaxed continuous states. Conversely, the diffusion process is extended to the discrete token space [4, 54, 51, 65], which removes the need for the aforementioned mapping via a specialized discrete diffusion formulation. In our work, we will focus on this latter perspective given its empirical potential as validated in recent work [51, 65, 23, 66]. Leveraging continuous-time Markov chain (CTMC) theory, the marginal distributions $p_{t}$ can be described by a family of linear ordinary differential equations
@@
To reverse this process, one may either model the density ratio $s_{\theta}(\mathbf{x})_{\mathbf{y}}\approx\frac{p_{t}(\mathbf{x})}{p_{t}(\mathbf{y})}$ given two sequences $\mathbf{x},\mathbf{y}\in\mathcal{X}\times\dots\times\mathcal{X}$ as in Uniform Score-Entropy Discrete Diffusion (SEDD) [51], or the denoised variate $\mathbf{x}_{\theta}(\mathbf{x}_{t},\alpha_{t})\approx\mathbf{x}$ directly as in masked diffusion model [65, 66]. In the former, the modeled density ratios induce a specialized reverse transition matrix $\bar{Q}_{t}$ that can be leveraged in Eq. 4. In the latter, $\mathbf{x}_{\theta}$ can be directly substituted for $\mathbf{x}$ in Eq. 5. In this work, we follow Sahoo et al. [65] that enforces zero-probability on the mask state $\mathbf{m}$ and keeps all un-masked state unchanged during reverse sampling. This induce a simplified (negative) variational lower bound under the continuous time limit
@@ 3 Method
We propose an end-to-end multi-modal diffusion model named Dual Diffusion Transformer (D-DiT) with a unified backbone that jointly models image and text distribution. More specifically, given image $\mathbf{x}^{(\text{img})}$ and text $\mathbf{x}^{(\text{txt})}$, we are interested at modeling the conditional distribution $p(\mathbf{x}^{(\text{img})}|\mathbf{x}^{(\text{txt})})$ and $p(\mathbf{x}^{(\text{txt})}|\mathbf{x}^{(\text{img})})$. The former is usually referred to as text-to-image generation and the latter is closely related to various image understanding tasks such as captioning and visual question answering.
@@ 3.1 Architecture
Inspired by the MM-DiT in SD3 [19], our proposed D-DiT is a Transformer-based model comprising two branches - one for processing image tokens and another for processing text tokens. The image and text tokens attend with each other in every attention layer. In D-DiT, the output of the image branch is the prediction of velocity defined in (2) with text conditioning, while the output for the text branch is the $\mathbf{x}^{(\text{txt})}$ prediction with image conditioning. The scalar timestep embedding modulates every layer’s feature map via AdaLN (adaptive layernorm) [56]. We only input the timestep $t$ to the model during image generation, as $\mathbf{x}^{(\text{txt})}_{t}$ has already implicitly carried the information of signal-to-noise ratio (the number of mask tokens in the sequence). In addition, we add a text encoder with bi-directional attention on top of the text branch of the diffusion model. Despite an asymmetric design of the model is not strictly required, having a text encoder on top of a DiT model allows us to easily adapt many existing text-to-to-image model such as SD3 and FLUX as pretrained backone for our D-DiT model. Note that the text encoder should not have causal mask as this will violate the masked diffusion process.
To reduce the computational cost associated with high-resolution images, we follow prior works on latent-space (image) diffusion [63] which compresses images from raw pixel space to a spatially compressed latent space obtained from a variational-autoencoder (VAE) trained with discriminator loss [18] and KL-divergence regularization [34].
@@ 3.2 Training
We propose a joint training objective for image-text joint modeling, which is essentially a joint denoising target that combines continuous and discrete diffusion. Formally, we use flow matching introduced in Section 2.1 to learn the conditional distribution of images and masked diffusion introduced in Section 2.2 to learn the conditional distribution of texts. During training, corrupted samples $\mathbf{x}_{t^{(\text{img})}},\mathbf{x}_{t^{(\text{txt})}}$  [^1] For better readability superscript $(\text{txt}),(\text{img})$ are omitted. are sampled from the corresponding forward corruption processes $q(\mathbf{x}_{t}|\mathbf{x})$ defined in (1) and (6) respectively and then we can calculate the diffusion loss for each modality
@@
where we use antithetic sampling for text diffusion timesteps $t_{i}$ by discretizing $(\delta,1]$ into $K$ points uniformly with $\delta$ being a small number to avoid numerical instability, for image diffusion we sample $t$ from the log-normal distribution. We do not corrupt the conditioning samples during training and as such the image diffusion timestep is always set to zero when predicting text distribution.
To sum up, the overall dual modality training loss is a simple weighted combination of the above single modality diffusion loss:
@@ 4.1 Experimental Setup
We implement our proposed framework based on the open-sourced SD3-medium model [19] [^2] https://huggingface.co/stabilityai/stable-diffusion-3-medium. We initialize the model weight of the DiT from the pretrained checkpoint and add a linear head on top of the text branch to do text denoising. Following SD3, we adopt the existing T5 encoder/tokenizer [61], and SD3’s image VAE, whose weights remain unchanged throughout all the experiments (except for the mask token embedding in T5). We remove the CLIP text encoders in the SD3 model due to its causal attention mask and for a simplified model structure. We use the special token <extra_id0> in T5’s vocabulary to represent the mask token in masked diffusion, as this token is used to mark the masked token in the mask pretraining process of original T5 model. In this way, we find the model can generate text reasonably well even without updating the weight of this token embedding. To further reduce the domain gap, we unfreeze the token embedding of <extra_id0> during the second stage of the training.
Different from multi-modal models that are built upon language models, our model has never been trained on text-only generation. In the preliminary experiments, we found that adding text-only target (i.e. un-conditional text generation) to the model does not influence its captioning performance significantly. An interesting future direction can be extending the proposed framework to model the marginal distribution of each modality.
@@
We train the model with three stages on publicly available datasets. The total image-text pairs we have used is roughly 40M. We list the detail of the dataset and training setup for each stage below, where all the training stages use the joint diffusion loss defined in (8).
@@
    Dual diffusion pretraining. The original SD3 model was only trained on ambient image-text pairs, and not solely on text data itself. To adapt D-DiT to text generation tasks, we train it on the joint diffusion loss for 60K iterations with a batch size of 512. The maximum text token length is truncated to 64 and we use image resolution of 256. The dataset used in this stage is re-captioned Datacomp-1b [21, 43] (the model has only seen around 30M images in this stage, which is less than $3\%$ of the total images in the dataset).
@@
    Continued pretraining on higher quality data. We then unfreeze the masked token embedding in T5 and train the model for 200k iterations on a image understanding dataset with rich textual description, which consists of the pretraining dataset from ShareGPT4V[11] (1.3M images) and OpenImages (1.9M subset with object detection annotations) [35] re-captioned by ShareCaptioner [^3] https://huggingface.co/Lin-Chen/ShareCaptioner. The text token length is set to 256 and image resolution is 256, with batch size of 512. Finetuning the mask token embedding reduces the domain gap as T5 encoder has not seen sequence filled with high percentage of mask token during its pretraining.
    However, as updating the mask token embedding requires backpropagating through the heavy T5 encoder, we freeze the mask embedding after this round of training on the image understanding dataset. We observe that the $\ell 2$ difference between the mask token embedding from different training iterations does not change much after 100k iterations.
    Moving forward, we conduct an optional high resolution model finetuning on the aforementioned image understanding dataset together with a newly added image generation dataset. Concretely, we collect a higher quality dataset with 10M images (9M re-captioned LAION-1024 and 1M midjourney images [^4] https://huggingface.co/datasets/CaptionEmporium/midjourney-niji-1m-llavanext). In this training stage, the image diffusion loss is calculated on the high quality image dataset whereas the text diffusion loss is calculated on the understanding dataset. We finetune the model for 80k iterations, with image resolution 512, text token length 256, and a batch size of 768. Only our 512$\times$512 model variant requires this training stage.
@@
    Visual instruction tuning. Finally, we finetune our model on an medley of instruction-tuning datasets to promote joint text-image conditioned text generation. We combine the LLaVA-Pretrain558K and LLaVA-v1.5-mix-665K visual instruction tuning datasets with the training splits for TextVQA and VizWiz and train for 25k iterations. Following the convention in LLaVA-1.5, the model is trained to distinguish between long-form answers and short answers, multiple choice answers, or captions via task-specific instruction prompts that come after the question such as ”Answer the question using a single word or phrase,” ”Answer with the option’s letter from the given choices directly,” or ”Describe the image concisely.”
@@ 4.2 Multi-modal Understanding
We first evaluate the visual understanding capabilities of D-DiT via the academic question answering benchmarks VQAv2 [26], VizWiz [8], OKVQA [53], GQA [30], POPE [44], as well as MME [20]. Due to the short-form nature of the questions, we perform sampling with 16 diffusion steps, and compare against a selection of multi-modal models, including I2T only and I2T + T2I models. Our results are summarized in Table 3. We note that our D-DiT as the only diffusion-only multi-modal model capable of visual question answering tasks, already boosts performance that is competitive with recent I2T + T2I models. Our model at 512 resolution outperforms Show-O on MME, GQA, and POPE, approaching performances of auto-regressive VLMs such as QWEN-VL and BLIP-2.
@@ 4.3 Text-to-image Generation
Besides the image-conditioned text generation, we also test model’s text-to-image generation capability. Following previous works, we evaluate our 512$\times$512 model after the second training stage on the GenEval benchmark, which measures model’s prompt following capability [25]. We follow the default setting in the open-sourced SD3 checkpoint where we use a Euler solver with 28 sampling steps and a CFG scale of $7.0$. We observe that the joint diffusion training does not cause catastrophic forgetting on the model, the fine-tuned D-DiT preserves the performance of original SD3 model and slightly improves on some of metrics such as colors after joint training. Qualitative evaluation samples are shown in Figure 3, where we observe the ability to generate highly aesthetic images is preserved.
@@ 4.4 Ablation Studies
As text-to-image diffusion models are trained on a large number of text-image pairs, one may raise the question, whether the representation learned throughout this process can be transferred to multi-modal understanding tasks? To answer this question, we perform an ablation study on the internal representation of a text-to-image diffusion model. We adapt several models into an image captioning model, including SD3, CLIP ViT L/14, and our D-DiT model. Among different internal layers in SD3, we find the feature from the 18th layer tend to perform the best in our preliminary experiment so it is used as the output feature. We add an GPT2 text decoder to the features extracted from SD3 and CLIP, and directly use D-DiT’s text output as results. We train all models with a mixture of recaptioned Datacomp, recaptioned OpenImages and captioning data from ShareGPT4V [11]. Concretely, we evaluate the quality of the captions generated from the models by asking GPT4 to do visual question answering according to the generated captions. The accuracy is listed in Table 4.
Similar to the trend observed in [75], directly using diffusion features as the prefix of a language decoder yield worse performance compared to language-supervised vision models like CLIP ViT [60]. Unfreezing the parameters of the diffusion backbone slightly improve the performance, but it still cannot match the CLIP encoder. This suggests that the representation from image diffusion models is not directly transferrable to the text embedding space where the decoder-only language model operates on. Instead of leveraging a separate language decoder, we use the text branch in the MM-DiT architecture to directly model the conditional text distribution, which notably boost the performance. This uncovers an intriguing property of MM-DiT model or potentially other bi-directional Transformers, that these models are good representation learners for estimating the likelihood of multi-modality data distribution.
We also conduct an ablation study with respect to the number of text diffusion sampling steps and study its influence on VQA accuracy with VQAv2 and captioning quality on COCO dataset. For VQAV2, which involves short text answers, good accuracy can be achieved with relatively few sampling steps. For the captioning task on MS-COCO, performance improves as the number of sampling steps increases, mirroring the trend observed by Sahoo et al. [65], where additional sampling steps lead to reduced perplexity.

### v2

@@ Abstract
Diffusion models have gained tremendous success in text-to-image generation, yet still struggle with visual understanding tasks, an area dominated by autoregressive vision-language models. We propose a large-scale and fully end-to-end diffusion model for multi-modal understanding and generation that significantly improves on existing diffusion-based multimodal models, and is the first of its kind to support the full suite of vision-language modeling capabilities. Inspired by the multimodal diffusion transformer (MM-DiT) and recent advances in discrete diffusion language modeling, we leverage a cross-modal maximum likelihood estimation framework that simultaneously trains the conditional likelihoods of both images and text jointly under a single loss function, which is back-propagated through both branches of the diffusion transformer. The resulting model is highly flexible and capable of a wide range of tasks including image generation, captioning, and visual question answering. Our model attained competitive performance compared to recent unified image understanding and generation models, demonstrating the potential of multimodal diffusion modeling as a promising alternative to autoregressive next-token prediction models.
@@ 2.1 Continuous Diffusion
with velocity field $\mathbf{v}(\mathbf{x}_{t},t)=\dot{\alpha_{t}}\mathbf{x}+\dot{\sigma_{t}}\mathbf{\epsilon}$. The ODE in (2) pushes the distribution of $\mathbf{x}_{t}$ from $p_{0}$ to $p_{T}$. To generate new samples, we can use neural networks to approximate $\mathbf{v}$ and then integrate ODE (2) backward in time starting from $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. A common choice of $\alpha_{t},\sigma_{t}$ in flow matching model is $\alpha_{t}=1-t,\sigma=t$ and therefore $\mathbf{v}=\epsilon-\mathbf{x}$, which corresponds to the optimal transport interpolant between two distribution $p_{0}$ and $p_{1}$ [48, 53]. The neural network for regressing the velocity field $\mathbf{v}$ in (2) is trained by optimizing the flow matching loss
@@ 2.2 Discrete Diffusion
In discrete diffusion, the variate $\mathbf{x}\in\mathcal{X}\times\dots\times\mathcal{X}$ has finite support over the product space of $\mathcal{X}=\{1,\dots,N\}$, where in language models $N$ is the vocabulary size of the token embedding. Generally, there are two ways to approach this modeling task. The first line of works [43, 12, 16, 55, 28] apply a continuous relaxation to the discrete variable and proceed with a continuous reformulation of the framework, allowing the application of the equations in Section 2.1. This greatly simplifies the diffusion modeling itself, but introduces a significant source of error in the mapping between discrete and relaxed continuous states. Conversely, the diffusion process is extended to the discrete token space [4, 57, 54, 68], which removes the need for the aforementioned mapping via a specialized discrete diffusion formulation. In our work, we will focus on this latter perspective given its empirical potential as validated in recent works [54, 68, 24, 69]. Leveraging continuous-time Markov chain (CTMC) theory, the marginal distributions $p_{t}$ can be described by a family of linear ordinary differential equations
@@
To reverse this process, one may either model the density ratio $s_{\theta}(\mathbf{x})_{\mathbf{y}}\approx\frac{p_{t}(\mathbf{x})}{p_{t}(\mathbf{y})}$ given two sequences $\mathbf{x},\mathbf{y}\in\mathcal{X}\times\dots\times\mathcal{X}$ as in [54], or the denoised variate $\mathbf{x}_{\theta}(\mathbf{x}_{t},\alpha_{t})\approx\mathbf{x}$ directly as in Sahoo et al. [68], Shi et al. [69]. In the former, the modeled density ratios induce a specialized reverse transition matrix $\bar{Q}_{t}$ that can be leveraged in Eq. 4. In the latter, $\mathbf{x}_{\theta}$ can be directly substituted for $\mathbf{x}$ in Eq. 5. In this work, we follow Sahoo et al. [68] that enforces zero-probability on the mask state $\mathbf{m}$ and keeps all un-masked state unchanged during reverse sampling. This induce a simplified (negative) variational lower bound under the continuous time limit
@@ 3 Method
We propose an end-to-end multi-modal diffusion model named Dual Diffusion Transformer (D-DiT) with a unified backbone that jointly models image and text distribution. More specifically, given image $\mathbf{x}^{(\text{img})}$ and text $\mathbf{x}^{(\text{txt})}$, we are interested in modeling the conditional distribution $p(\mathbf{x}^{(\text{img})}|\mathbf{x}^{(\text{txt})})$ and $p(\mathbf{x}^{(\text{txt})}|\mathbf{x}^{(\text{img})})$. The former is usually referred to as text-to-image generation and the latter forms the basis for various image understanding tasks such as captioning and visual question answering.
@@ 3.1 Architecture
Inspired by the MM-DiT in SD3 [20], our proposed D-DiT is a Transformer-based model comprising two branches - one for processing image tokens and another for processing text tokens. The image and text tokens attend to each other in every attention layer. In D-DiT, the output of the image branch is the prediction of velocity defined in (2) with text conditioning, while the output for the text branch is the $\mathbf{x}^{(\text{txt})}$ prediction with image conditioning. The scalar timestep embedding modulates every layer’s feature map via AdaLN (adaptive layernorm) [59]. We only input the timestep information $t$ to the model for image generation, as $\mathbf{x}^{(\text{txt})}_{t}$ implicitly encodes this information as the ratio of masked tokens in the sequence. In addition, we add a text encoder with bi-directional attention on top of the text branch of the diffusion model. While the asymmetry between image and text branches is not strictly required, having a text encoder on top of a DiT model allows us to easily adapt many existing text-to-to-image models such as SD3 and FLUX as pretrained backbones for our D-DiT model (see Table 9 in the Appendix for a comparison). Note that the text encoder should not use a causal mask as this will violate the masked diffusion process.
To reduce the computational cost associated with modeling high-resolution images, we follow prior works on latent-space (image) diffusion [66], which encode images from the raw pixel space into a spatially compressed latent space obtained from a variational autoencoder (VAE) trained with a discriminator loss [19] and KL-divergence regularization [36].
@@ 3.2 Training
We propose a combined training objective for image-text modeling, which is essentially a joint denoising target that combines continuous and discrete diffusion. Formally, we use flow matching introduced in Section 2.1 to learn the conditional distribution of images and masked diffusion introduced in Section 2.2 to learn the conditional distribution of texts. During training, corrupted samples $\mathbf{x}_{t^{(\text{img})}},\mathbf{x}_{t^{(\text{txt})}}$  [^1] For better readability superscript $(\text{txt}),(\text{img})$ are omitted. are drawn from the corresponding forward corruption processes $q(\mathbf{x}_{t}|\mathbf{x})$ defined in (1) and (6) respectively. We then calculate the diffusion loss for each modality as
@@
In text diffusion, we use antithetic sampling [35] for timesteps $t_{i}$ by discretizing $(\delta,1]$ into $K$ points uniformly with $\delta$ being a small number to avoid numerical instability. In image diffusion, we sample $t$ from the log-normal distribution. We do not corrupt the conditioning samples during training, i.e., the image diffusion timestep is always set to zero when predicting text distribution and vice versa.
In summary, the overall dual modality training loss is a simple weighted combination of the above single modality diffusion loss:
@@ 4.1 Experimental Setup
We implement our proposed framework based on the open-sourced SD3-medium model [20] [^2] https://huggingface.co/stabilityai/stable-diffusion-3-medium. We initialize the model weights of the DiT from the pretrained checkpoint and add a linear head on top of the text branch for text denoising. Following SD3, we adopt the existing T5 encoder/tokenizer [64], and SD3’s image VAE, whose weights remain unchanged throughout all the experiments (except for the mask token embedding in T5). We remove the CLIP text encoders in the SD3 model due to its causal attention mask and for a simplified model structure. We use the special token <extra_id0> in T5’s vocabulary to represent the mask token in masked diffusion, as this token is used to mark the masked token in the mask pretraining process of original T5 model. This way, we find the model can generate text reasonably well even without updating the weight of this token embedding. To further reduce the domain gap, we unfreeze the token embedding of <extra_id0> during the second stage of the training.
Different from multi-modal models that are built upon language models, our model has never been trained on text-only generation. In preliminary experiments, we found that adding a text-only target (i.e. unconditional text generation) to the model does not influence its captioning performance significantly. An interesting future direction can be extending the proposed framework to model the marginal distribution of each modality.
@@
We train the model in three stages on publicly available datasets. The total number of image-text pairs used is roughly 40M. We list the details of the dataset and training setup for each stage below, where all the training stages use the joint diffusion loss defined in (8).
@@
    Dual diffusion pretraining. The original SD3 model was only trained on ambient image-text pairs, and not solely on text data itself. To adapt D-DiT to text generation tasks, we train it on the joint diffusion loss for 60K iterations with a batch size of 512. The maximum text token length is truncated to 64 and we use an image resolution of 256. The dataset used in this stage is re-captioned Datacomp-1b [22, 45] (the model has only seen around 30M images in this stage, which is less than $3\%$ of the total images in the dataset).
@@
    Continued pretraining on higher quality data. We then unfreeze the masked token embedding in T5 and train the model for 200k iterations on an image understanding dataset with rich textual description, which consists of the pretraining dataset from ShareGPT4V[11] (1.3M images) and OpenImages (1.9M subset with object detection annotations) [37] re-captioned by ShareCaptioner [^3] https://huggingface.co/Lin-Chen/ShareCaptioner. The text token length is set to 256 and image resolution to 256, with a batch size of 512. Finetuning the mask token embedding reduces the domain gap as T5 encoder has not seen sequences filled with a high percentage of mask tokens during its pretraining.
    However, as updating the mask token embedding requires backpropagating through the large T5 encoder, we freeze the mask embedding after this round of training on the image understanding dataset. We observe that the $\ell 2$ difference between the mask token embedding from different training iterations does not change much after 100k iterations.
    Here, we may conduct an optional high resolution model finetuning on the aforementioned image understanding dataset together with a higher quality dataset with 10M images (9M re-captioned LAION-1024 and 1M midjourney images [^4] https://huggingface.co/datasets/CaptionEmporium/midjourney-niji-1m-llavanext). In this training stage, the image diffusion loss is calculated on the high quality image dataset whereas the text diffusion loss is calculated on the understanding dataset. We finetune the model for 80k iterations, with image resolution 512, text token length 256, and a batch size of 768. Only our 512$\times$512 model variant requires this training stage.
@@
    Visual instruction tuning. Finally, we finetune our model on a medley of instruction-tuning datasets to promote joint text-image conditioned text generation. We combine the LLaVA-Pretrain558K and LLaVA-v1.5-mix-665K visual instruction tuning datasets with the training splits for TextVQA and VizWiz and train for 50k iterations. Following the convention in LLaVA-1.5, the model is trained to distinguish between long-form and short answers, multiple choice answers, or captions via task-specific instruction prompts that come after the question, e.g. "Answer the question using a single word or phrase," or "Describe the image concisely."
@@ 4.2 Multi-modal Understanding
We first evaluate the visual understanding capabilities of D-DiT via the academic question answering benchmarks VQAv2 [27], VizWiz [8], OKVQA [56], GQA [32], POPE [46], as well as MME [21]. Due to the short-form nature of the questions, we perform sampling with 16 diffusion steps, and compare against a selection of multi-modal models, including I2T only and I2T + T2I models. Our results are summarized in Table 3. We note that our D-DiT is the only diffusion-only multi-modal model capable of visual question answering tasks, already boosting performance that is competitive with recent I2T + T2I models. Our model at 512 resolution outperforms Show-O on MME, GQA, and POPE, approaching the performance of auto-regressive VLMs such as QWEN-VL and BLIP-2.
@@ 4.3 Text-to-image Generation
Besides the image-conditioned text generation, we also test model’s text-to-image generation capability. Following previous works, we evaluate our 512$\times$512 model after the second training stage on the GenEval benchmark, which measures model’s prompt following capability [26]. We follow the default setting in the open-sourced SD3 checkpoint where we use a Euler solver with 28 sampling steps and a CFG scale of $7.0$. We observe that the joint diffusion training does not cause catastrophic forgetting on the model; the fine-tuned D-DiT preserves the performance of the original SD3 model and slightly improves on some metrics such as color accuracy after joint training. Qualitative evaluation samples are shown in Figure 3, where we observe that the ability to generate highly aesthetic images is preserved.
@@ 4.4 Ablation Studies
As text-to-image diffusion models are trained on a large number of text-image pairs, one may raise the question of whether the representation learned throughout this process can be transferred to multi-modal understanding tasks. To answer this question, we perform an ablation study on the internal representation of a text-to-image diffusion model. We adapt several models into an image captioning model, including SD3, CLIP ViT L/14, and our D-DiT model. Among different internal layers in SD3, we find the feature from the 18th layer tends to perform the best in our preliminary experiments so it is used as the output feature. We add a GPT2 text decoder to the features extracted from SD3 and CLIP, and directly use D-DiT’s text output as results. We train all models with a mixture of recaptioned Datacomp, recaptioned OpenImages and captioning data from ShareGPT4V [11]. Concretely, we evaluate the quality of the captions generated from the models by asking GPT4 to do visual question answering according to the generated captions. The accuracy is listed in Table 4.
Similar to the trend observed in [78], directly using diffusion features as the prefix of a language decoder yields worse performance compared to language-supervised vision models like CLIP ViT [63]. Unfreezing the parameters of the diffusion backbone slightly improves the performance, but it still cannot match the performance of the CLIP encoder. This suggests that the representation from image diffusion models is not directly transferable to the text embedding space where the decoder-only language model operates on. Instead of leveraging a separate language decoder, we use the text branch in the MM-DiT architecture to directly model the conditional text distribution, which notably boosts the performance. This uncovers an intriguing property of MM-DiT models, and potentially other bi-directional Transformers: that these models are good representation learners for estimating the likelihood of multi-modality data distributions.
We also conduct an ablation study with respect to the number of text diffusion sampling steps and study its influence on VQA accuracy with VQAv2 and captioning quality on the COCO dataset. For VQAv2, which involves short text answers, good accuracy can be achieved with relatively few sampling steps. For the captioning task on MS-COCO, performance improves as the number of sampling steps increases, mirroring the trend observed by Sahoo et al. [68], where additional sampling steps lead to reduced perplexity.